OpenStudy (anonymous):

What type of variation is modled in the table? A.) direct B.) joint C.) combined D.) inverse

OpenStudy (anonymous):

x= 2 4 6 Y= 24 12 8

OpenStudy (anonymous):

its not a because 12*2=24 4*3=12

OpenStudy (anonymous):

Okay i get that.

OpenStudy (anonymous):

and it isnt joint because 12*2*24 does not equal 4*3*12

OpenStudy (anonymous):

I get that as well. so would it be inverse?

OpenStudy (anonymous):

it is inverse.

OpenStudy (anonymous):

I mean combined

OpenStudy (anonymous):

Because x=k/y

OpenStudy (anonymous):

How would that be?

OpenStudy (anonymous):

Bye

OpenStudy (anonymous):

You sure?